OpenStudy (anonymous):

two angles of a triangle have measures of 50 degrees and 65 degrees. which of the following could not be a measure of an exterior angle of the triangle?

OpenStudy (compassionate):

Well, the angles are 50 and 65. A triangle adds up to 180 degrees. So, 50 + 50 + 15 = 115 - 180 = 65. Really, your third angle is 65; and the exterior angles can't be greater or lesser than this

OpenStudy (compassionate):

From Google: An exterior angle of a triangle is equal to the sum of the opposite interior angles. For more on this see Triangle external angle theorem. If the equivalent angle is taken at each vertex, the exterior angles always add to 360° In fact, this is true for any convex polygon, not just triangles. So, if we take two opposite angles of a triangle. We know our angles are 50, 65, 65. This is a special triangle. so add the sum of the opposite angles to get your possible solutions. For example 50 + 65 yields 115 for the straight line that passes on points 50 and 65